OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (−6, 3). y = three halvesx − 1 y = three halvesx + 3 y = −two thirdsx − 1 y = −two thirdsx + 3

OpenStudy (anonymous):

@gypsy1274 This is the second to last problem, I think I got it, but I am not to sure.

OpenStudy (debbieg):

what do you think the answer is?

OpenStudy (anonymous):

So we, would start out with y-y1 = m (x-x1) right? and I am not sure.

OpenStudy (anonymous):

Yes, but remember what a perpendicular slope is...

OpenStudy (anonymous):

Is it set up like y-3 = 3/2 (x-(-6))

OpenStudy (debbieg):

That would give you a line parallel to the given line. What is the relationship between the slope of the line given in the problem (3/2) and the slope of a line that is perpendicular to it?

OpenStudy (anonymous):

Y + 3 = 3/2 (x- 6) ?

OpenStudy (anonymous):

Your first equation was closer. It is the slope....

OpenStudy (anonymous):

What do you know about a perpendicular slope? Draw a set of perpendicular lines on a graph and find their slopes.

OpenStudy (anonymous):

Oh! Do I have to flip the fraction?

OpenStudy (anonymous):

Perpendicular slope = negative reciprocal

OpenStudy (anonymous):

Oh so - 2/3

OpenStudy (anonymous):

Now, you got it!!

OpenStudy (anonymous):

Ok! let me try this out!

OpenStudy (anonymous):

I got; Y= -2/3-1

OpenStudy (anonymous):

So did I. Great job!!

OpenStudy (anonymous):

Wooo! I have one more problem, which Is a word problem.

OpenStudy (anonymous):

Be sure to pick the x back up off the floor!!

OpenStudy (anonymous):

Explain, in complete sentences, the relationship between the equations of parallel lines.

OpenStudy (anonymous):

Think about the slope.

OpenStudy (anonymous):

You distribute the slope in both of the problems?

OpenStudy (anonymous):

Think about the relationship of the slopes of parallel lines.

OpenStudy (anonymous):

\(y=6x+4\) is parallel to \(y=6x-7\)

OpenStudy (anonymous):

Ohh, so they are parallel to each other? and you can graph them on the same chart?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Do you feel like you understand?

OpenStudy (anonymous):

I do :)

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

Thank you! I got 100% on that test!

OpenStudy (anonymous):

Great. If you can, try to do some more similar problems tomorrow to make sure you remember and understand.

OpenStudy (anonymous):

OK, thank you :)

OpenStudy (anonymous):

Your welcome.